# Inverse Virtual Try-On: Generating Multi-Category Product-Style Images from Clothed Individuals

### v1

@@ Abstract
While virtual try-on (VTON) systems aim to render a garment onto a target person image, this paper tackles the novel task of virtual try-off (VTOFF), which addresses the inverse problem: generating standardized product images of garments from real-world photos of clothed individuals. Unlike VTON, which must resolve diverse pose and style variations, VTOFF benefits from a consistent and well-defined output format – typically a flat, lay-down-style representation of the garment – making it a promising tool for data generation and dataset enhancement. However, existing VTOFF approaches face two major limitations: (i) difficulty in disentangling garment features from occlusions and complex poses, often leading to visual artifacts, and (ii) restricted applicability to single-category garments (e.g., upper-body clothes only), limiting generalization. To address these challenges, we present Text-Enhanced MUlti-category Virtual Try-Off (TEMU-VTOFF), a novel architecture featuring a dual DiT-based backbone with a modified multimodal attention mechanism for robust garment feature extraction. Our architecture is designed to receive garment information from multiple modalities like images, text, and masks to work in a multi-category setting. Finally, we propose an additional alignment module to further refine the generated visual details. Experiments on VITON-HD and Dress Code datasets show that TEMU-VTOFF sets a new state-of-the-art on the VTOFF task, significantly improving both visual quality and fidelity to the target garments. Our project page is available at https://temu-vtoff-page.github.io/.
@@ 1 Introduction
Despite the recent success of GANs goodfellow2014GAN and Latent Diffusion Models (LDMs) rombach2022ldm in image translation tasks siarohin2019first ; ren2023pi ; isola2017image ; tumanyan2023plug , current VTOFF solutions face notable limitations. Existing models velioglu2024tryoffdiff ; xarchakos2024tryoffanyone struggle to accurately reconstruct catalog images from dressed human inputs. This limitation arises from a fundamental architectural mismatch – these approaches repurpose VTON pipelines by merely reversing the input-output roles, without addressing the unique challenges of the VTOFF task. Moreover, the high visual variability of real-world images – due to garment wear category (e.g., upper-body), pose changes, and occlusions – makes it difficult for these models to robustly extract garment features while preserving fine-grained patterns. On the opposite side, we design a dedicated architecture tailored for the VTOFF task, accounting for the unique challenges of complex positions and occlusions in person images as opposed to a flat catalog garment.
Recent advances in diffusion models demonstrate that DiT-based architectures peebles2023scalable , especially when combined with flow-matching objectives lipman2022flow , surpass traditional U-Net and DDPM-based approaches rombach2022ldm . Inspired by these findings, we propose TEMU-VTOFF, a Text-Enhanced MUlti-category Virtual Try-Off architecture based on a dual-DiT framework. Specifically, we exploit the representational strength of DiT in two distinct ways: (i) the first Transformer component focuses on extracting fine-grained garment features from complex, detail-rich person images; and (ii) the second DiT is specialized for generating the clean, in-shop version of the garment. To support this design, we further adapt the base DiT architecture to accommodate the task-specific input modalities. To further enhance alignment, we introduce an external garment aligner module and a novel supervision loss that leverages clean garment references as guidance, further improving quality of generated images.
@@
  • Multimodal Hybrid Attention. We introduce a novel attention mechanism that integrates garment textual descriptions into the generative process by linking them with person-specific features. This helps the model synthesize occluded or ambiguous garment regions more accurately.
@@ 2 Related Work
Virtual Try-On. As one of the most popular tasks within the fashion domain, VTON has been widely studied over the past decades by the computer vision and graphics communities due to its interesting research challenges and the practical potential bai2022single ; cui2021dressing ; fele2022c ; ren2023cloth ; ge2021disentangled . Existing methods are broadly categorized into warping-based chen2023size ; xie2023gpvton ; yan2023linking and warping-free approaches zhu2023tryondiffusion ; morelli2023ladi ; baldrati2023multimodal ; zeng2024cat ; chong2025catvton , with a growing shift from GAN-based goodfellow2020generative to diffusion-based ho2020denoising ; song2020denoising frameworks for better image fidelity and stability. Warping-based methods usually follow a two-stage pipeline: first warping the garment using TPS bookstein1989principal , flow, or landmarks to align it with the body, then fusing it with the person image. VITON han2018viton , CP-VTON wang2018toward , and their variants improve garment alignment and synthesis quality, but often produce artifacts due to imperfect warping. To mitigate this, warping-free methods leverage diffusion models to bypass explicit deformation zhu2023tryondiffusion ; morelli2023ladi ; xu2024ootdiffusion ; choi2024idmimproving employing modified cross-attention or self-attention to directly condition generation on garment features, often using CLIP-based encoders radford2021learning . However, these pre-trained encoders tend to lose fine-grained texture details, prompting methods like StableVITON kim2023stableviton to introduce dedicated garment encoders and attention mechanisms, albeit at a higher computational cost. Lately, DiT-based works jiang2024fitdit show the benefits of Transformer-based diffusion models for high-fidelity garment to person transfer. While most existing works focus on generating dressed images from separate garment and person inputs, the inverse problem (i.e., reconstructing clean, standalone garment representations from worn images) remains underexplored.
@@
Virtual Try-Off. While VTON has been extensively studied for synthesizing images of a person wearing a target garment, the recently proposed VTOFF task shifts the focus toward garment-centric reconstruction, aiming to extract a clean, standardized image of a garment worn by a person. TryOffDiff velioglu2024tryoffdiff introduces this task by leveraging a diffusion-based model with SigLIP zhai2023sigmoid conditioning to recover high-fidelity garment images. Building on this direction, TryOffAnyone xarchakos2024tryoffanyone addresses the generation of tiled garment images from dressed photos for applications like outfit composition and retrieval. By integrating garment-specific masks and simplifying the Stable Diffusion pipeline through selective Transformer tuning, it achieves both quality and efficiency. In both cases, these works have been designed for single-category scenarios, thus limiting their potential application to generate wider, more diverse data collections. On a different line, Any2AnyTryon guo2025any2anytryonleveragingadaptiveposition is not a native VTOFF method, but it leverages a LoRA-based module hu2022lora to fine-tune FLUX flux2024 for this task. Though these works collectively reflect a growing shift from person-centric synthesis to garment-centric understanding, there are still limitations like frequent garment structural artifacts (e.g., in shape, neckline, waist) and on colors and textures of generated outputs. We hypothesize that this mismatch is due to a too generic architectural choice, untailored for the specific needs of the VTOFF setting. In this work, we focus on existing VTOFF open problems, such as multi-category adaptation, occlusions, and complex human poses, and propose a novel VTOFF-specific architecture enhanced with text and fine-grained mask conditioning and optimized with a garment aligner component that can improve the quality of generated garments.
@@ 3 Methodology
Figure 2: Overview of our method. The feature extractor $F_{E}$ processes spatial inputs (noise, masked image, binary mask), and global inputs (model image via AdaLN). The intermediate keys and values $\bm{K}^{l}_{\text{extractor}}$,$\bm{V}^{l}_{\text{extractor}}$ are injected into the corresponding hybrid blocks of the garment generator $F_{D}$. Then, the main DiT model generates the final garment leveraging the proposed MHA module. We align our model with a diffusion loss for the noise estimate and an alignment loss with clean, DINOv2 features of the target garment.
Preliminary. The latest diffusion models are a family of generative architectures that work by corrupting a ground-truth image $\bm{z}_{0}$ following a flow-matching schedule lipman2022flow defined as
@@
Then, a diffusion model estimates back the injected noise $\epsilon_{t}$ through a Diffusion Transformer (DiT) peebles2023scalable , obtaining a prediction $\hat{\bm{z}_{0}}$. In Stable Diffusion 3 (SD3) esser2024scaling , the 16-channel latent $\bm{z}_{t}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 16}$ is obtained projecting the original RGB image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ with a variational autoencoder $\mathcal{E}$ kingma2013vae , obtaining $\bm{z}=\mathcal{E}(\bm{x})$, with $H,W$ being height and width of the image, and $f=8$ the spatial compression ratio of the autoencoder. Finally, the model is trained according to an MSE loss function $\mathcal{L}_{\text{diff}}$:
@@
Overview. An overview of our method is shown in Fig. 2. The purpose of this framework is to generate an in-shop version of the garment worn by the person. The critical design choice is to process correctly the dressed person image in order to extract meaningful information to be injected in the denoising process. For this reason, we encompass a dual-DiT architecture, based on SD3, whose models are deputed to two different purposes. Firstly, we design the first DiT as a feature extractor $F_{E}$ that encodes the model image $\bm{x}_{\text{model}}$ and outputs its intermediate layer features at timestep $t=0$ and not from subsequent timesteps, as we are interested in extracting clean features from $F_{E}$. This block is trained with a diffusion loss to generate the person image. Once trained, this model outputs meaningful key and value features of the dressed person. Secondly, the main DiT generates the garment $\bm{x}_{g}$ leveraging the intermediate features from $F_{E}$ in a modified textual-enhanced attention module. The two DiT architectures can be manipulated along three axes: (i) the Transformer projector before the DiT to account for additional inputs; (ii) the Modulation Space (of every DiT block) to account for different input modalities (such as text and images); (iii) the Attention operator (in all layers) for fine-grained conditioning.
@@ 3.1 DiT Feature Extractor
We design the feature extractor $F_{E}$ as a DiT, working with two different input types: a global input with the person image $\bm{x}_{model}\in\mathbb{R}^{H\times W\times 3}$ leveraged by the modulation layers of $F_{E}$, and a local spatial input as the channel-wise concatenation $\bm{z}_{t}^{\prime}=[\bm{z}_{t},M,\bm{x}_{M}]\in\mathbb{R}^{h\times w\times 33}$ of the latent $\bm{z}_{t}$, the encoded latent of the masked person image $\bm{x}_{M}=\mathcal{E}(I_{M})\in\mathbb{R}^{h\times w\times 16}$ and the interpolated binary mask $M\in\mathbb{R}^{h\times w\times 1}$ encoded through the Transformer projector $\mathcal{P}:\mathbb{R}^{h\times w\times 33}\rightarrow\mathbb{R}^{S\times d}$, with $S$ as sequence length and $d$ as embedding dimension. We train it as a single DiT separated from the garment generator $F_{D}$. We extract the keys and values $\bm{K}^{l}_{\text{extractor}},\bm{V}^{l}_{\text{extractor}}$ from every attention block $A^{l}$ of $F_{E}$, with $l=1,\ldots,N$, as shown in Fig. 2, because they bring valuable visual information about the garment features in the person image.
Mask Conditioning with modified Transformer Projector. We investigate whether fine-grained clothing attributes – such as garment shape, edge refinement, and neckline structure – can be inferred from text alone. These attributes are inherently structural and spatial, which poses a challenge for language-based conditioning due to its inherently diffuse nature. We posit that a segmentation mask can act as a “hard” discriminator, providing precise spatial constraints that complement the “soft” and distributed cues present in text.
To this end, we propose to use the mask $M$ and the associated masked person image $\bm{\bm{x}}_{M}$ as an additional conditioning of the extractor $F_{E}$ and we get a new latent $\bm{z}_{t}^{\prime}=[\bm{z}_{t},M,\bm{x}_{M}]\in\mathbb{R}^{h\times w\times 33}$. We modify the original SD3 DiT projector $\mathcal{P_{\text{orig}}}:\mathbb{R}^{h\times w\times 16}\rightarrow\mathbb{R}^{S\times d}$ by adding zero-initialized convolutional layers on the channel dimension, thus accounting for the increased channel-size of our input $\bm{z}_{t}^{\prime}$ and obtaining the final projector $\mathcal{P}$. We train the module $F_{E}$ alone, detached from the dual DiT $F_{D}$, according to the diffusion loss $L_{\text{diff}}$ defined as follows:
where we align the generated features with the person image, thus encouraging $F_{E}$ attention keys and values to retain valuable information of the garment worn by the model.
Visual-only Modulation Space and Attention. The modulation space of the feature extractor $F_{E}$ follows AdaLN huang2017arbitrary and receives features encoded with CLIP ViT-L radford2021learning and Open-CLIP bigG/14 cherti2023reproducible . For simplicity, we will refer to this operation as $\operatorname{CLIP}(c)$, with $c$ being a textual or visual input. The feature extractor $F_{E}$ does not necessitate text as input, because we want to capture the garment details from $\bm{x}_{\text{model}}$, without “steering” the output process with text. For this reason, we encode the visual projection $\bm{e}^{v}_{\text{pool}}=\operatorname{CLIP}(\bm{x}_{\text{model}})\in\mathbb{R}^{2048}$, which is subsequently used to modulate the latent $\bm{z}_{t}$ from AdaLN estimated scale $\gamma$ and shift $\beta$ as follows:
$\bm{y}_{t}$ $=\operatorname{MLP}(t,\bm{e}^{v}_{\text{pool}})$
$\bm{z}_{t}$ $\leftarrow\gamma(\bm{y}_{t})\bm{z}_{t}+\beta(\bm{y}_{t})$ (4)
using an MLP to jointly encode the pooled vector $\bm{e}^{v}_{\text{pool}}$ together with the timestep $t\in(0,1)$.
From Eq. 4 we notice how AdaLN modulation can shift the distribution $\bm{z}_{t}$ according to “appearance” or “style” of our conditioning pooled input. This justifies our architectural choice, as the latent $\bm{z}_{t}$ in $F_{E}$ inherits the garment information from $\bm{e}^{v}_{\text{pooled}}$ through AdaLN and then propagates it through its layers. Following the reasoning above, we stick to visual-only self-attention modules for $F_{E}$ DiT blocks with queries, keys and values as $\bm{Q}=\bm{Q}_{\bm{z}_{t}},\bm{K}=\bm{K}_{\bm{z}_{t}},\bm{V}=\bm{V}_{\bm{z}_{t}}$. Although this design strategy proves effective for $F_{E}$, we show in Sec. 3.2 that our main DiT $F_{D}$ needs to incorporate three diverse types of input, leading to a new Multimodal Hybrid Attention (MHA) module.
3.2 Text-Enhanced Garment Try-Off
Previous VTON and VTOFF works leverage visual-only inputs, as they claim to have a better transfer of information from the conditioning input to the DiT generation stream. While this approach turns out to be effective for VTON tasks, it has two key shortcomings in the VTOFF setting. First, it prevents our method from generating occluded garments (e.g., a bodysuit worn under a pair of pants). Second, it limits the application of VTOFF to single-category data collections (e.g., upper-body only). Instead, we propose a new, text-enhanced DiT $F_{D}$ to solve these two issues.

### v2

@@ Abstract
Virtual try-on (VTON) has been widely explored for rendering garments onto person images, while its inverse task, virtual try-off (VTOFF), remains largely overlooked. VTOFF aims to recover standardized product images of garments directly from photos of clothed individuals. This capability is of great practical importance for e-commerce platforms, large-scale dataset curation, and the training of foundation models. Unlike VTON, which must handle diverse poses and styles, VTOFF naturally benefits from a consistent output format in the form of flat garment images. However, existing methods face two major limitations: (i) exclusive reliance on visual cues from a single photo often leads to ambiguity, and (ii) generated images usually suffer from loss of fine details, limiting their real-world applicability. To address these challenges, we introduce TEMU-VTOFF, a Text-Enhanced MUlti-category framework for VTOFF. Our architecture is built on a dual DiT-based backbone equipped with a multimodal attention mechanism that jointly exploits image, text, and mask information to resolve visual ambiguities and enable robust feature learning across garment categories. To explicitly mitigate detail degradation, we further design an alignment module that refines garment structures and textures, ensuring high-quality outputs. Extensive experiments on VITON-HD and Dress Code show that TEMU-VTOFF achieves new state-of-the-art performance, substantially improving both visual realism and consistency with target garments. Code and models are available at: https://temu-vtoff-page.github.io/.
@@ 1 Introduction
Despite the success of GANs (Goodfellow et al., 2014) and Latent Diffusion Models (LDMs) (Rombach et al., 2022) in image translation tasks (Siarohin et al., 2019; Ren et al., 2023b; Isola et al., 2017; Tumanyan et al., 2023), current VTOFF solutions face notable limitations. Existing models (Velioglu et al., 2024; Xarchakos and Koukopoulos, 2024) struggle to accurately reconstruct catalog images from dressed human inputs. This limitation arises from a fundamental architectural mismatch: these approaches repurpose VTON pipelines by merely reversing the input-output roles, without addressing the unique challenges of the VTOFF task. Moreover, the high visual variability of real-world images – due to garment wear category (e.g., upper-body), pose changes, and occlusions – makes it difficult for these models to robustly extract garment features while preserving fine-grained patterns. On the opposite side, we design a dedicated architecture tailored for the VTOFF task.
Recent advances in diffusion models demonstrate that DiT-based architectures (Peebles and Xie, 2023), especially when combined with flow-matching objectives (Lipman et al., 2023), surpass traditional U-Net and DDPM-based approaches (Rombach et al., 2022). Inspired by these findings, we propose TEMU-VTOFF, a Text-Enhanced MUlti-category Virtual Try-OFF architecture based on a dual-DiT framework. Specifically, we exploit the representational strength of DiT in two distinct ways: (i) the first Transformer component focuses on extracting fine-grained garment features from complex, detail-rich person images; and (ii) the second DiT is specialized for generating the clean, in-shop version of the garment. To support this design, we further adapt the base DiT architecture to accommodate the task-specific input modalities. To further enhance alignment, we introduce an external garment aligner module and a novel supervision loss that leverages clean garment references as guidance, further improving quality of generated images.
@@
  • Multimodal Hybrid Attention. We introduce a novel attention mechanism that integrates garment textual descriptions into the generative process by linking them with person-specific features. This helps the dual-DiT architecture synthesize the garments more accurately.
@@ 2 Related Work
Virtual Try-On. As a core task in the fashion domain, VTON has been extensively studied by the computer vision and graphics communities due to its practical potential (Bai et al., 2022; Cui et al., 2021; Fele et al., 2022; Ren et al., 2023a). Existing methods are broadly categorized into warping-based (Chen et al., 2023; Xie et al., 2023; Yan et al., 2023) and warping-free approaches (Zhu et al., 2023; Morelli et al., 2023; Baldrati et al., 2023; Zeng et al., 2024; Chong et al., 2025), with a growing shift from GAN-based (Goodfellow et al., 2020) to diffusion-based frameworks (Ho et al., 2020; Song et al., 2021). VITON (Han et al., 2018) and its variants (Wang et al., 2018; Choi et al., 2021; Kang et al., 2021) improve garment alignment and synthesis quality, but often produce artifacts due to imperfect warping. To mitigate this, warping-free methods leverage diffusion models to bypass explicit deformation (Zhu et al., 2023; Morelli et al., 2023; Xu et al., 2025; Choi et al., 2024), using modified cross- or self-attention to condition generation on garment features. However, these pre-trained encoders tend to lose fine-grained texture details, motivating methods like StableVITON (Kim et al., 2024) to add dedicated garment encoders and attention modules, at increased computational cost. Lately, DiT-based works (Jiang et al., 2025; Zhu et al., 2024) show the benefits of Transformer-based diffusion models for high-fidelity garment to person transfer. Finally, some models adopt more elaborate conditioning strategies. For instance, LOTS introduces a pair-former module for handling multiple inputs (Girella et al., 2025), while LEFFA learns a flow field from averaged cross-attention maps and employs learnable tokens to stabilize attention values (Zhou et al., 2025). While most works focus on generating dressed images from separate person and garment inputs, the inverse problem (i.e., reconstructing clean garment representations from worn images) remains underexplored.
@@
Virtual Try-Off. While VTON has been extensively studied for synthesizing images of a person wearing a target garment, the recently proposed VTOFF task shifts the focus toward garment-centric reconstruction, aiming to extract a clean, standardized image of a garment worn by a person. TryOffDiff (Velioglu et al., 2024) introduces this task by leveraging a diffusion-based model with SigLIP (Zhai et al., 2023) conditioning to recover high-fidelity garment images. Building on this direction, TryOffAnyone (Xarchakos and Koukopoulos, 2024) addresses the generation of tiled garment images from dressed photos for applications like outfit composition and retrieval. By integrating garment-specific masks and simplifying the Stable Diffusion pipeline through selective Transformer tuning, it balances quality and efficiency. In both cases, these works have been designed for single-category scenarios, limiting scalability to diverse data collections. Recent efforts have started to address these limitations. MGT (Velioglu et al., 2025) extends VTOFF to multi-category scenarios by incorporating class-specific embeddings to handle diverse clothing types within a unified model. More ambitious efforts unify VTON and VTOFF in a single framework: Voost (Lee and Kwak, 2025) proposes a single DiT to learn both tasks, while One Model For All (Liu et al., 2025) introduces a partial diffusion mechanism to achieve a similar goal. On a different line, Any2AnyTryon (Guo et al., 2025) adapts FLUX (Labs, 2024) using a LoRA-based module (Hu et al., 2022) for this task. While these works reflect a shift from person-centric synthesis to garment-centric understanding, they still suffer from structural artifacts (e.g., shape, neckline, waist) and inaccurate colors or textures. We hypothesize that this mismatch is due to a too generic architectural choice, not tailored for the specific needs of the VTOFF setting. In this work, we focus on existing VTOFF open problems, such as multi-category adaptation, occlusions, and complex human poses, and propose a novel VTOFF-specific architecture enhanced with text and fine-grained mask conditioning and optimized with a garment aligner component that can improve the quality of generated garments.
Conditioning Methods in Diffusion Models. To overcome the limitations of text-only conditioning, many schemes leverage additional visual inputs such as segmentation maps, bounding boxes, poses, and points (Sun et al., 2024; Li et al., 2023; Chen et al., 2024; Nie et al., 2024; Lin et al., 2024; Wang et al., 2024). Prominent methods like ControlNet (Zhang et al., 2023) and T2I-Adapter (Mou and others, 2024) inject spatial conditions via auxiliary networks, while IP-Adapter (Ye et al., 2023) uses separate attention branches more suited to U-Nets than DiTs. Other works focus on unifying multiple conditions, either through modular controllers like Uni-ControlNet (Zhao et al., 2023) and dedicated adapters (Lin et al., 2025), or by concatenating visual embeddings directly into the Transformer input sequence (Tan et al., 2025; Wang et al., 2025; Xiao et al., 2025). Although these methods are effective for general personalization tasks (i.e., placing an object from one image into another), they lack the fine-grained conditioning mechanism necessary to extract specific garment data from person images, a gap we address to unlock the VTOFF task.
@@ 3 Methodology
Figure 2: Overview of our method. The feature extractor $F_{E}$ processes spatial inputs (noise, masked image, binary mask) and global inputs (model image via AdaLN). The intermediate keys and values $\bm{K}^{l}_{\text{extractor}}$, $\bm{V}^{l}_{\text{extractor}}$ are injected into the corresponding hybrid blocks of the garment generator $F_{D}$. The main DiT then produces the final garment using the proposed MHA module. Training combines a diffusion loss on noise prediction with an alignment loss with DINOv2 features of the target garment.
Preliminaries. The latest diffusion models are a family of generative architectures that corrupt a ground-truth image $\bm{z}_{0}$ according to a flow-matching schedule (Lipman et al., 2023) defined as
@@
Then, a diffusion model estimates back the injected noise $\epsilon_{t}$ through a Diffusion Transformer (DiT) (Peebles and Xie, 2023), obtaining a prediction $\hat{\bm{z}_{0}}$. In Stable Diffusion 3 (SD3) (Esser et al., 2024), the 16-channel latent $\bm{z}_{t}\in\mathbb{R}^{\frac{H}{8}\times\frac{W}{8}\times 16}$ is obtained projecting the original RGB image $\bm{x}\in\mathbb{R}^{H\times W\times 3}$ with a variational autoencoder $\mathcal{E}$ (Kingma and Welling, 2013), obtaining $\bm{z}=\mathcal{E}(\bm{x})$, with $H,W$ being height and width of the image, and $f=8$ the spatial compression ratio of the autoencoder. Finally, the model is trained according to an MSE loss function $\mathcal{L}_{\text{diff}}$:
@@
Overview. An overview of our method is shown in Fig. 2. The objective is to generate an in-shop version of the garment worn by the person. A critical design choice lies in processing the dressed person image so as to extract meaningful information for injection into the denoising process. To this end, we adopt a dual-DiT architecture, built upon SD3, with the two models assigned to complementary roles. Firstly, we design the first DiT as a feature extractor $F_{E}$ that encodes the model image $\bm{x}_{\text{model}}$ and outputs its intermediate layer features at timestep $t=0$ and not from subsequent timesteps, as we are interested in extracting clean features from $F_{E}$. This block is trained with a diffusion loss to generate the person image. Once trained, this model outputs meaningful key and value features of the dressed person. Secondly, the main DiT generates the garment $\bm{x}_{g}$ leveraging the intermediate features from $F_{E}$ in a modified textual-enhanced attention module.
@@ 3.1 DiT Feature Extractor
At inference time, the only available input is the clothed person image $\bm{x}_{\text{model}}\in\mathbb{R}^{H\times W\times 3}$, from which we also extract the mask. To encode this information, we compute the visual projection
$\bm{e}^{v}_{\text{pool}}=\operatorname{CLIP}(\bm{x}_{\text{model}})\in\mathbb{R}^{2048},$
which is then used to modulate the latent $\bm{z}_{t}$ through the AdaLN-estimated scale $\gamma$ and shift $\beta$:
$\bm{y}_{t}$ $=\operatorname{MLP}(t,\bm{e}^{v}_{\text{pool}}),$
$\bm{z}_{t}$ $\leftarrow\gamma(\bm{y}_{t})\bm{z}_{t}+\beta(\bm{y}_{t}).$ (3)
Existing VTON approaches rely on two visual inputs: the target garment and the person. In our case, however, the model can rely only on person features, from which it is more complex to extract the garment features. This shortcoming makes the CLIP vector $\bm{e}^{v}_{\text{pool}}$ the real bottleneck in a unified architecture setting, as the CLIP projection alone is too coarse to properly encode this information.
To address this, we propose introducing a dedicated feature extractor $F_{E}$, allowing $F_{D}$ to concentrate exclusively on the garment generation task. The architecture of $F_{E}$ mirrors that of the main SD3 DiT module $F_{D}$, with the only difference being its input layer, which is adapted to handle additional visual inputs in the channel dimension. The inputs are a global input with the person image $\bm{x}_{model}\in\mathbb{R}^{H\times W\times 3}$ encoded as $\bm{e}^{v}_{\text{pool}}$ and leveraged by the modulation layers of $F_{E}$, and a local spatial input as the channel-wise concatenation $\bm{z}_{t}^{\prime}=[\bm{z}_{t},M,\bm{x}_{M}]\in\mathbb{R}^{h\times w\times 33}$ of the latent $\bm{z}_{t}$, the encoded latent of the masked person image $x_{M}=\mathcal{E}(x_{model}\odot M)\in\mathbb{R}^{h\times w\times 16}$, where $\mathcal{E}$ is the SD3 VAE encoder, $h=H/f$ and $w=W/f$ denote the spatial dimensions after downsampling by the factor $f=8$, and the interpolated binary mask $M\in\mathbb{R}^{h\times w\times 1}$ encoded through the Transformer projector $\mathcal{P}:\mathbb{R}^{h\times w\times 33}\rightarrow\mathbb{R}^{S\times d}$, with $S$ as sequence length and $d$ as embedding dimension.
This design choice is central to our method, as each layer output of the feature extractor $F_{E}$ retains meaningful intermediate representations of both the person and the garment. Leveraging these features offers three key advantages: (i) instead of the collapsed CLIP representation, we obtain expanded features of dimension $S\times d$; (ii) the $L$ layers of $F_{E}$ capture information at multiple granularities, progressing from coarse to fine (Avrahami et al., 2025; Skorokhodov et al., 2025), so that each layer $l$ conveys a different level of detail about the same image; (iii) since $F_{E}$ shares the same architecture as $F_{D}$, the features extracted at layer $l$ from $F_{E}$ are naturally better aligned with those of $F_{D}$. Motivated by these considerations, we extract the keys $\bm{K}^{l}_{\text{extractor}}$ and values $\bm{V}^{l}_{\text{extractor}}$ from every layer $l$ of $F_{E}$.
3.2 Dual-DiT Text-Enhanced Garment Try-Off
Without loss of generality, we will omit the index $l$ when referring to $\bm{Q}^{l}$ $\bm{K}^{l}$ and $\bm{V}^{l}$ of $F_{E}$ and $F_{D}$, since the conditioning scheme is applied uniformly across all layers. Given the extracted features $\bm{K}_{\text{extractor}}$ and $\bm{V}_{\text{extractor}}$, we propose to modify the SD3 attention scheme to incorporate such information, leading to our Multimodal Hybrid Attention (MHA).